# An Objective Improvement Approach

### v3

@@ 5. Choosing Better Strategies
    All local improvements can be applied. A strategy $\sigma^{\prime}$ is a local improvement to a strategy $\sigma$ if $f_{\sigma^{\prime}}(\mathsf{val})<f_{\sigma}(\mathsf{val})$ holds for the current valuation $\mathsf{val}$ (Section 5.1).
@@
    If the current valuation $\mathsf{val}$ does not define a pair of strategies $\sigma$ for both players and has no local improvements, then a better strategy $\sigma^{\prime}$ can be found applying only switches from and to edges that already have offset 00 (Section 5.2).
@@ 5.1. Local Improvements
The simplest and most common case of creating better strategies $\sigma^{\prime}$ from a valuation for the objective $f_{\sigma}$ for a strategy $\sigma$ is to consider local improvements. Much like local improvements in strategy iteration approaches, local improvements consider, for each vertex $v$, a successor $v^{\prime}\neq\sigma(v)$, such that $\mathsf{offset}(\mathsf{val},(v,v^{\prime}))<\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$ for the current valuation $\mathsf{val}$, which is optimal for the objective function $f_{\sigma}$.
@@
If $\mathsf{val}$ is an optimal valuation for the linear programming problem at Line 5 of Algorithm 3 and $f_{\sigma^{\prime}}(\mathsf{val})<f_{\sigma}(\mathsf{val})$, then $\sigma^{\prime}$ is better than $\sigma$.
@@
The valuation $\mathsf{val}$ is, being an optimal solution for the objective $f_{\sigma}$, a solution to the system of inequations $H$. For a solution $\mathsf{val}^{\prime}$ that is optimal for $f_{\sigma^{\prime}}$, we thus have $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})\leq f_{\sigma^{\prime}}(\mathsf{val})<f_{\sigma}(\mathsf{val})$, which implies that $\sigma^{\prime}$ is better than $\sigma$ according to Definition 4.1 of a better strategy.
@@ 5.2. No Local Improvements
We define for a valuation $\mathsf{val}$ optimal for $f_{\sigma}$ (like the $\mathsf{val}$ produced in Line 5 of Algorithm 3):
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be a DPG, $\sigma$ a strategy for both players, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 3 for $f_{\sigma}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. If $\mathsf{val}$ does not define strategies of both players, then there is a better strategy $\sigma^{\prime}$ such that, for all $v\in V$, $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$.
@@
When using the most natural choice, $E_{\mathsf{val}}^{\sigma}=E_{\mathsf{val}}\cup\{(v,\sigma(v))\mid v\in V\}$, this translates into keeping all transitions, for which the offset is not 00, while changing some of those for which the offset already is 00. This is a slightly surprising choice, since to progress one has to improve on the transitions whose offset is positive, and ignore those with offset 0.
@@ 5.3. Games with Efficient Objective Improvement
The simplest condition for efficiently finding better strategies is the existence of local improvements. In particular, it is easy to find, for a given valuation $\mathsf{val}$, strategies $\sigma^{\prime}$ for both players such that $f_{\sigma^{\prime}}(\mathsf{val})\leq f_{\sigma^{\prime\prime}}(\mathsf{val})$ holds for all strategies $\sigma^{\prime\prime}$. When there are local improvements, we can obtain a better strategy simply by applying them. This leaves the case in which there are no local improvements, but where $\mathsf{val}$ also does not define strategies for the two players. We have seen that we can obtain a better strategy by only swapping edges, for which the inequations are sharp (Lemma 4).
@@
To do this efficiently, it helps if there are always $|V|$ inequations that are sharp for a valuation $\mathsf{val}$ that minimises $f_{\sigma}(\mathsf{val})$ for some $\sigma$. Of course, if $\mathsf{val}$ is a solution returned by the simplex method, then there must be at least $|V|$ sharp inequations, as $\mathsf{val}$ is the solution defined by making $|V|$ inequations sharp (namely those that form the basis). By requiring that there are exactly $|V|$ sharp inequations we require that an optimal valuation defines a basis. We call such a set of inequations $H$ and games that define them sharp games.
@@
(Sharp Game) A game $\mathcal{G}$ is called sharp for a valuation $\mathsf{val}$ if, and only if, $\mathsf{val}$ satisfies exactly $|V|$ inequations $H$ computed by $\mathsf{Inequations}(\mathcal{G})$ sharply. It is called sharp if no solution satisfies strictly more than $|V|$ inequations sharply.
@@
The second condition, which allows us to identify better strategies efficiently, is to assume that, for every strategy $\sigma$ for both players, if a valuation $\mathsf{val}$ defined by a basis is not optimal for $f_{\sigma}$ under the constraints $H$, then there is a single basis change that improves it. We call such games improving. Being improving is the property we really need, while sharpness is a property that allows us to almost surely make a game improving by applying small noise on the objective functions.
@@
Bases define valuations, and valuations define (sets of strategies): as we have used in Lemma 3, it is easy to identify for a valuation $\mathsf{val}$ the pair of strategies $\sigma$ (or the pairs of strategies if there are many) for a whom $f_{\sigma}(\mathsf{val})$ is smallest among all pairs of strategies; they are simply those pairs that minimise the offset for every vertex.
We therefore say that a value defines these pairs of strategies, and that a basis defines the pairs of strategies defined by the valuation it defines.
@@
(Improving Game) A game $\mathcal{G}$ is called $\sigma$-improving for an objective function $f_{\sigma}$ defined by a joined strategy $\sigma$ and a non-optimal valuation $\mathsf{val}$ defined by a basis for $f_{\sigma}$ if, and only if, there is a single basis change that leads to a better strategy (having a smaller value of $f_{\sigma}$) defined by the new basis.
A game $\mathcal{G}$ is called improving if it is improving for all objective functions $f_{\sigma}$ defined by a joined strategy $\sigma$ and all valuations $\mathsf{val}$ defined by bases that are not optimal w.r.t. $f_{\sigma}$.
We call a valuation $\mathsf{val}^{\prime}$ whose basis can be obtained from that of $\mathsf{val}$ by a single change to the basis of $\mathsf{val}$ a neighbouring valuation to $\mathsf{val}$. We show that, for improving games, we can refine the result of Lemma 4 so that the better strategy $\sigma^{\prime}$ also guarantees $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<f_{\sigma}(\mathsf{val})$ for some neighbouring valuation $\mathsf{val}^{\prime}$ to $\mathsf{val}$.
This allows us to consider $O(|E|)$ basis changes and, where they define a valuation, to seek optimal strategies for a given valuation. Finding an optimal strategy for a given valuation is straightforward.
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be an improving DPG, $\sigma$ a strategy for both players that is not co-optimal, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 3 for $f_{\sigma}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. Then there is (at least) one neighbouring valuation $\mathsf{val}^{\prime\prime}$ to $\mathsf{val}$ such that there is a better strategy $\sigma^{\prime}$ that satisfies $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})<f_{\sigma}(\mathsf{val})$.
@@
By Lemma 4, we can obtain a better strategy $\sigma^{\prime}$ from transitions in $E_{\mathsf{val}}^{\sigma}$; for $\sigma^{\prime}$ therefore $f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})$ holds, and $\mathsf{val}$ is not optimal for $f_{\sigma}$.
Let $\mathbf{b}$ be a basis that defines $\mathsf{val}$. As $\mathcal{G}$ is improving, it is $\sigma^{\prime}$-improving, so that there is basis $\mathbf{b}^{\prime}$ that neighbours $\mathbf{b}$ and defines a valuation $\mathsf{val}^{\prime}$ with $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})$.
@@
While we can restrict the selection of $\sigma^{\prime}$ to the strategies that comply with the restriction $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$, there is no particular reason for doing so; as soon as we have a neighbouring valuation $\mathsf{val}^{\prime}$, we can identify a pair of strategies $\sigma^{\prime}$ for which $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})$ is minimal and select $\sigma^{\prime}$ if $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<f_{\sigma}(\mathsf{val})$ holds.
@@
As an alternative to considering all neighbouring bases of $\mathbf{b}$, determining their valuation, and determining the objective function for $\sigma^{\prime}$, we can also determine the reduced cost for exchanging a basis based on the best improvement obtained by any function derived from $E^{\sigma}_{\mathsf{val}}$ and use this to locate an improvement.
@@ 5.4. Making Games Sharp and Improving
Before defining the gap of the game, we recall that $\mathsf{val}_{\sigma}$ is the value of the joint strategy $\sigma$ (whereas $\mathsf{val}$ denotes the outcome of the optimisation problem). We use this valuation to argue that a small distortion of the edge weights cannot turn non-co-optimal strategies into co-optimal ones, and we define the gap of the game to reason about how small is small enough to retain non-co-optimality. The value of a joint strategy itself is useful for this because it is much easier to access than the result of optimisation. As a valuation of a non-co-optimal joint strategy is not a solution of a game, we will encounter negative offsets: they occur for exactly those edges, for which the associated inequation is not satisfied by $\mathsf{val}_{\sigma}$.
@@
In a strategy improvement approach, where there is an edge $e$ with $-\mathsf{offset}(\mathsf{val}_{\sigma},e)>0$, this would constitute a profitable switch, allowing the player to improve their return. While we do not use it this way, we argue that the gap $\gamma$ allows us to infer that, when we change all edge weights by a sufficiently small factor (which we can derive from $\gamma_{\sigma}$, or $\gamma$, and the contraction of the game), a non-co-optimal strategy $\sigma$ is still non-co-optimal after adjusting the weights this way.
@@
If we assume that $\sigma$ is not co-optimal for $\mathcal{G}$, then we have an edge $e=(v,v^{\prime})$ with $-\mathsf{offset}(\mathsf{val}_{\sigma},e)=\gamma_{\sigma}\geq\gamma$. Using $\mathsf{offset}^{\prime}$ to indicate the use of $w_{e}^{\prime}$ for $\mathcal{G}^{\prime}$, by applying the triangle inequality we get
$\begin{array}[]{cl}|\mathsf{offset}^{\prime}(\mathsf{val}_{\sigma}^{\prime},e)-\mathsf{offset}(\mathsf{val}_{\sigma},e)|\!\!\!\!&\leq|\mathsf{val}_{\sigma}^{\prime}(v)-\mathsf{val}_{\sigma}(v)|+|w_{e}-w_{e}^{\prime}|+\lambda_{e}|\mathsf{val}_{\sigma}(v^{\prime})-\mathsf{val}_{\sigma}^{\prime}(v^{\prime})|\\ &<\frac{\gamma}{2}+\frac{1-\lambda^{*}}{2}\gamma+\lambda_{e}\frac{\gamma}{2}\leq\gamma.\\ \end{array}$
Together with the fact that $-\mathsf{offset}(\mathsf{val}_{\sigma},e)\geq\gamma$, this provides $\mathsf{offset}^{\prime}(\mathsf{val}_{\sigma}^{\prime},e)<0$.
@@
Note that these distributions are continuous. They could be replaced by any other distribution over these intervals that has weight 00 for all individual points.
@@
What we show is that it happens with probability 00 that $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ are bases and define the same solution in $\mathcal{G}$. If $\mathbf{b}_{1}$ or $\mathbf{b}_{2}$ is not a basis (which would happen if the inequations are not linearly independent, e.g. if a vertex $v$ does not appear at all in these inequations) we are done, too.
@@
For the valuation $\mathsf{val}_{1}$ defined by $\mathbf{b}_{1}$ to be equal to the valuation $\mathsf{val}_{2}$ defined by $\mathbf{b}_{2}$, the weight for the edge $e$ (after adding the drawn distortion) needs to be exactly $w_{e}^{\prime}$. (This is a necessary condition, but not necessarily a sufficient condition.) There is at most one value for the disturbance that would provide for this, and this disturbance for the weight of $e$ is sampled with a likelihood of 00.
@@
Starting in $v$, $\sigma$ defines a run $\rho=e_{1}e_{2}e_{3}\ldots$ in the form of a “lasso path”, which consists of a (possibly empty) initial path $e_{1},\ldots,e_{k}$, followed by an infinitely often repeated cycle $e_{1}^{\prime},\ldots,e_{\ell}^{\prime}$, where the only vertex that occurs twice on the path $e_{1},\ldots,e_{k},e_{1}^{\prime},\ldots,e_{\ell}^{\prime}$ is the vertex reached before $e_{1}^{\prime}$ and after $e_{\ell}^{\prime}$, while all other vertices occur only once.
@@
This leaves the contribution of the denominator of the fraction from the second form, which is the nominator of the term $1-\prod_{j=1}^{\ell}\lambda_{e_{j}^{\prime}}$. But since the term is between 00 and $1$, its value is strictly smaller than the value of its denominator; we thus have
$\mathsf{nom}(1-\prod_{j=1}^{\ell}\lambda_{e_{j}^{\prime}})<\prod_{j=1}^{\ell}\mathsf{denom}(\lambda_{e_{j}^{\prime}})\ .$
To obtain $-\mathsf{offset}(\mathsf{val}_{\sigma},e)$, we have to compare with a run starting in $v$, but taking the edge $e\neq(v,\sigma(v))$ first.
This defines a run $\rho^{\prime}=d_{1}d_{2}d_{3}\ldots$ (with $d_{1}=e$), again in the form of a “lasso path” that consists of an initial path $d_{1},\ldots,d_{k^{\prime}}$, followed by an infinitely often repeated cycle $d_{1}^{\prime},\ldots,d_{\ell^{\prime}}^{\prime}$, where the only vertex that occurs twice on the path $d_{1},\ldots,d_{k},d_{1}^{\prime},\ldots,d_{\ell^{\prime}}^{\prime}$ is the vertex reached before $d_{1}^{\prime}$ and after $d_{\ell^{\prime}}^{\prime}$, while all other vertices occur only once.
@@
This again leaves the contribution of the denominator of the fraction from the second form, which is the nominator of the term $1-\prod_{j=1}^{{\ell^{\prime}}}\lambda_{d_{j}^{\prime}}$. Again a term between 00 and $1$, its value is strictly smaller than the value of its denominator; we thus have
$\mathsf{nom}(1-\prod_{j=1}^{{\ell^{\prime}}}\lambda_{d_{j}^{\prime}})<\prod_{j=1}^{\ell^{\prime}}\mathsf{denom}(\lambda_{d_{j}^{\prime}})\ .$
@@
Obviously, the nominator will differ by at least one, so that $\gamma$ is at least the reciprocal of the largest such value.
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be an improving DPG for a given set of positive numbers $\{\alpha_{e}>0\mid e\in E\}$, $\sigma$ a strategy for both players, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 3 for the adjusted function $f_{\sigma}^{\prime}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. Then there is a neighbouring valuation $\mathsf{val}^{\prime\prime}$ to $\mathsf{val}$ such that there is a better strategy $\sigma^{\prime}$ that satisfies $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})<f_{\sigma}(\mathsf{val})$.

### v4

@@ 5. Choosing Better Strategies
    All local improvements can be applied. A strategy $\sigma^{\prime}$ is a local improvement to a strategy $\sigma$ if $f_{\sigma^{\prime}}(\mathsf{val})<f_{\sigma}(\mathsf{val})$ holds for the current solution $\mathsf{val}$ (Section 5.1).
@@
    If the current solution $\mathsf{val}$ does not define a pair of strategies $\sigma$ for both players and has no local improvements, then a better strategy $\sigma^{\prime}$ can be found applying only switches from and to edges that already have offset 0 (Section 5.2).
@@ 5.1. Local Improvements
The simplest and most common case of creating better strategies $\sigma^{\prime}$ from a solution for the objective $f_{\sigma}$ for a strategy $\sigma$ is to consider local improvements. Much like local improvements in strategy iteration approaches, local improvements consider, for each vertex $v$, a successor $v^{\prime}\neq\sigma(v)$, such that $\mathsf{offset}(\mathsf{val},(v,v^{\prime}))<\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$ for the current solution $\mathsf{val}$, which is optimal for the objective function $f_{\sigma}$.
@@
If $\mathsf{val}$ is an optimal solution for the linear programming problem at Line 5 of Algorithm 2 and $f_{\sigma^{\prime}}(\mathsf{val})<f_{\sigma}(\mathsf{val})$, then $\sigma^{\prime}$ is better than $\sigma$.
@@
The solution $\mathsf{val}$ is, being an optimal solution for the objective $f_{\sigma}$, a solution to the system of inequations $H$. For a solution $\mathsf{val}^{\prime}$ that is optimal for $f_{\sigma^{\prime}}$, we thus have $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})\leq f_{\sigma^{\prime}}(\mathsf{val})<f_{\sigma}(\mathsf{val})$, which implies that $\sigma^{\prime}$ is better than $\sigma$ according to Definition 4.1 of a better strategy.
This allows us to identify local improvements cheaply.
Corollary 4.
If $\mathsf{val}$ is an optimal solution for the linear programming problem at Line 5 of Algorithm 2, there is an edge $(v,v^{\prime})$ such that $\mathsf{offset}(\mathsf{val},(v,v^{\prime}))<\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$, and $\sigma^{\prime}$ is a strategy for both players s.t. $\mathsf{offset}(\mathsf{val},(v,v^{\prime}))\geq\mathsf{offset}(\mathsf{val},(v,\sigma^{\prime}(v)))$ holds for all edges $(v,v^{\prime})$, then $\sigma^{\prime}$ is better than $\sigma$.
We say that $\mathsf{val}$ identifies these strategies.
@@ 5.2. No Local Improvements
We define for a solution $\mathsf{val}$ optimal for $f_{\sigma}$ (like the $\mathsf{val}$ produced in Line 5 of Algorithm 2):
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be a DPG, $\sigma$ a strategy for both players, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 2 for $f_{\sigma}$. If $\mathsf{val}$ does not define strategies of both players, then there is a better strategy $\sigma^{\prime}$ such that, for all $v\in V$, $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$.
@@
When using the most natural choice, $E_{\mathsf{val}}^{\sigma}=E_{\mathsf{val}}\cup\{(v,\sigma(v))\mid v\in V\}$, this translates into keeping all transitions, for which the offset is not 0, while changing some of those for which the offset already is 0. This is a slightly surprising choice, since to progress one has to improve on the transitions whose offset is positive, and ignore those with offset 0.
While Lemma 5 also holds when there is a local improvement, we would not use it then, as finding local improvements is much easier.
@@ 5.3. Games with Efficient Objective Improvement
The simplest condition for efficiently finding better strategies is the existence of local improvements. In particular, it is easy to find, for a given solution $\mathsf{val}$, strategies $\sigma^{\prime}$ for both players such that $f_{\sigma^{\prime}}(\mathsf{val})\leq f_{\sigma^{\prime\prime}}(\mathsf{val})$ holds for all strategies $\sigma^{\prime\prime}$. When there are local improvements, we can obtain a better strategy simply by applying them. This leaves the case in which there are no local improvements, but where $\mathsf{val}$ also does not define strategies for the two players. We have seen that we can obtain a better strategy by only swapping edges, for which the inequations are sharp (Lemma 5).
@@
To do this efficiently, it helps if there are always $|V|$ inequations that are sharp for a solution $\mathsf{val}$ that minimises $f_{\sigma}(\mathsf{val})$ for some $\sigma$. Of course, if $\mathsf{val}$ is a solution returned by the simplex method, then there must be at least $|V|$ sharp inequations, as $\mathsf{val}$ is the solution defined by making $|V|$ inequations sharp (namely those that form the basis). By requiring that there are exactly $|V|$ sharp inequations we require that an optimal solution defines a basis. We call such a set of inequations $H$ and games that define them sharp games.
@@
(Sharp Game) A game $\mathcal{G}$ is called sharp for a solution $\mathsf{val}$ if, and only if, $\mathsf{val}$ satisfies exactly $|V|$ inequations $H$ computed by $\mathsf{Inequations}(\mathcal{G})$ sharply. It is called sharp if no solution satisfies strictly more than $|V|$ inequations sharply.
@@
The second condition, which allows us to identify better strategies efficiently, is to assume that, for every strategy $\sigma$ for both players, if a solution $\mathsf{val}$ defined by a basis is not optimal for $f_{\sigma}$ under the constraints $H$, then there is a single basis change that improves it. We call such games improving. Being improving is the property we really need, while sharpness is a property that allows us to almost surely make a game improving by applying small noise on the objective functions.
@@
Bases define solutions, and solutions identify (sets of) strategies: as we have used in Corollary 4, it is easy to identify for a solution $\mathsf{val}$ the pair of strategies $\sigma$ (or the pairs of strategies if there are many) for whom $f_{\sigma}(\mathsf{val})$ is the smallest among all pairs of strategies; they are simply those pairs that minimise the offset for every vertex.
We therefore say that a solution identifies these pairs of strategies, and that a basis defines the pairs of strategies defined by the solution it defines.
@@
(Improving Game) A game $\mathcal{G}$ is called $\sigma$-improving for an objective function $f_{\sigma}$ defined by a joined strategy $\sigma$ and a non-optimal solution $\mathsf{val}$ defined by a basis for $f_{\sigma}$ if, and only if, there is a single basis change that leads to a better strategy (having a smaller value of $f_{\sigma}$) defined by the new basis.
A game $\mathcal{G}$ is called improving if it is improving for all objective functions $f_{\sigma}$ defined by a joined strategy $\sigma$ and all solutions $\mathsf{val}$ defined by bases that are not optimal w.r.t. $f_{\sigma}$.
We call a solution $\mathsf{val}^{\prime}$ whose basis can be obtained from that of $\mathsf{val}$ by a single change to the basis of $\mathsf{val}$ a neighbouring solution to $\mathsf{val}$. We show that, for improving games, we can refine the result of Lemma 5 so that the better strategy $\sigma^{\prime}$ also guarantees $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<f_{\sigma}(\mathsf{val})$ for some neighbouring solution $\mathsf{val}^{\prime}$ to $\mathsf{val}$.
This allows us to consider $O(|E|)$ basis changes and, where they define a solution, to seek optimal strategies for a given solution. Finding an optimal strategy for a given solution is straightforward.
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be an improving DPG, $\sigma$ a strategy for both players that is not co-optimal, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 2 for $f_{\sigma}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. Then there is (at least) one neighbouring solution $\mathsf{val}^{\prime\prime}$ to $\mathsf{val}$ such that there is a better strategy $\sigma^{\prime}$ that satisfies $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})<f_{\sigma}(\mathsf{val})$.
@@
By Lemma 5, we can obtain a better strategy $\sigma^{\prime}$ from transitions in $E_{\mathsf{val}}^{\sigma}$; for $\sigma^{\prime}$ therefore $f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})$ holds, and $\mathsf{val}$ is not optimal for $f_{\sigma^{\prime}}$.
Let $\mathbf{b}$ be a basis that defines $\mathsf{val}$. As $\mathcal{G}$ is improving, it is $\sigma^{\prime}$-improving, so that there is a basis $\mathbf{b}^{\prime}$ that neighbours $\mathbf{b}$ and defines a solution $\mathsf{val}^{\prime}$ with $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})$.
@@
While we can restrict the selection of $\sigma^{\prime}$ to the strategies that comply with the restriction $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$, there is no particular reason for doing so; as soon as we have a neighbouring solution $\mathsf{val}^{\prime}$, we can identify a pair of strategies $\sigma^{\prime}$ for which $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})$ is minimal and select $\sigma^{\prime}$ if $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<f_{\sigma}(\mathsf{val})$ holds.
@@
As an alternative to considering all neighbouring bases of $\mathbf{b}$, determining their valuation (and checking that they are solutions), and determining the objective function for $\sigma^{\prime}$, we can also determine the reduced cost for exchanging a basis based on the best improvement obtained by any function derived from $E^{\sigma}_{\mathsf{val}}$ and use this to locate an improvement.
@@ 5.4. Making Games Sharp and Improving
Before defining the gap of the game, we recall that $\mathsf{val}_{\sigma}$ is the value of the joint strategy $\sigma$ (whereas $\mathsf{val}$ denotes the outcome of the optimisation problem). We use this solution to argue that a small distortion of the edge weights cannot turn non-co-optimal strategies into co-optimal ones, and we define the gap of the game to reason about how small is small enough to retain non-co-optimality. The value of a joint strategy itself is useful for this because it is much easier to access than the result of optimisation. As a valuation of a non-co-optimal joint strategy is not a solution of a game, we will encounter negative offsets: they occur for exactly those edges, for which the associated inequation is not satisfied by $\mathsf{val}_{\sigma}$.
@@
In a strategy improvement approach, where there is an edge $e$ with $\mathsf{offset}(\mathsf{val}_{\sigma},e)>0$, this would constitute a profitable switch, allowing the player to improve their return. While we do not use it this way, we argue that the gap $\gamma$ allows us to infer that, when we change all edge weights by a sufficiently small factor (which we can derive from $\gamma_{\sigma}$, or $\gamma$, and the contraction of the game), a non-co-optimal strategy $\sigma$ is still non-co-optimal after adjusting the weights this way.
@@
If we assume that $\sigma$ is not co-optimal for $\mathcal{G}$, then we have an edge $e=(v,v^{\prime})$ with $\mathsf{offset}(\mathsf{val}_{\sigma},e)=\gamma_{\sigma}\geq\gamma$. Using $\mathsf{offset}^{\prime}$ to indicate the use of $w_{e}^{\prime}$ for $\mathcal{G}^{\prime}$, by applying the triangle inequality we get
[…]
Together with the fact that $\mathsf{offset}(\mathsf{val}_{\sigma},e)\geq\gamma$, this provides $\mathsf{offset}^{\prime}(\mathsf{val}_{\sigma}^{\prime},e)<0$.
@@
Note that these distributions are continuous. They could be replaced by any other distribution over these intervals that has weight 0 for all individual points.
@@
What we show is that it happens with probability 0 that $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ are bases and define the same valuations (regardless of whether or not these valuations are solutions) in $\mathcal{G}$. If $\mathbf{b}_{1}$ or $\mathbf{b}_{2}$ is not a basis (which would happen if the inequations are not linearly independent, e.g. if a vertex $v$ does not appear at all in these inequations) we are done, too.
@@
For the valuation $\mathsf{val}_{1}$ defined by $\mathbf{b}_{1}$ to be equal to the valuation $\mathsf{val}_{2}$ defined by $\mathbf{b}_{2}$, the weight for the edge $e$ (after adding the drawn distortion) needs to be exactly $w_{e}^{\prime}$. (This is a necessary condition, but not necessarily a sufficient condition.) There is at most one value for the disturbance that would provide for this, and this disturbance for the weight of $e$ is sampled with a likelihood of 0.
@@
Starting in $v$, $\sigma$ defines a run $\rho=e_{1}e_{2}e_{3}\ldots$ in the form of a “lasso path”, which consists of a (possibly empty) initial path $e_{1},\ldots,e_{k}$, followed by an infinitely often repeated cycle $e_{1}^{\prime},\ldots,e_{\ell}^{\prime}$, where the only edge that occurs twice on the path $e_{1},\ldots,e_{k},e_{1}^{\prime},\ldots,e_{\ell}^{\prime}$ is the edge reached before $e_{1}^{\prime}$ and after $e_{\ell}^{\prime}$, while all other edges occur only once.
@@
This leaves the contribution of the denominator of the fraction from the second form, which is the nominator of the term $1-\prod_{j=1}^{\ell}\lambda_{e_{j}^{\prime}}$. But since the term is between 0 and $1$, its value is strictly smaller than the value of its denominator; we thus have for
$n_{1}=\mathsf{nom}(1-\prod_{j=1}^{\ell}\lambda_{e_{j}^{\prime}})<\prod_{j=1}^{\ell}\mathsf{denom}(\lambda_{e_{j}^{\prime}})\leq\prod\limits_{v\in V}\mathsf{denom}(\lambda_{(v,\sigma(v))})$
that
$\mathsf{common^{\prime}}\cdot n_{1}\mbox{ is a denominator for }\mathsf{out}(\rho)\ .$
To obtain $\mathsf{offset}(\mathsf{val}_{\sigma},e)$, we have to compare with a run starting in $v$, but taking the edge $e\neq(v,\sigma(v))$ first.
This defines a run $\rho^{\prime}=d_{1}d_{2}d_{3}\ldots$ (with $d_{1}=e$), again in the form of a “lasso path” that consists of an initial path $d_{1},\ldots,d_{k^{\prime}}$, followed by an infinitely often repeated cycle $d_{1}^{\prime},\ldots,d_{\ell^{\prime}}^{\prime}$, where the only edge that occurs twice on the path $d_{1},\ldots,d_{k^{\prime}},d_{1}^{\prime},\ldots,d_{\ell^{\prime}}^{\prime}$ is the edge reached before $d_{1}^{\prime}$ and after $d_{\ell^{\prime}}^{\prime}$, while all other edges occur only once.
@@
This again leaves the contribution of the denominator of the fraction from the second form, which is the nominator of the term $1-\prod_{j=1}^{{\ell^{\prime}}}\lambda_{d_{j}^{\prime}}$. Again a term between 0 and $1$, its value is strictly smaller than the value of its denominator; we thus have for
$n_{2}=\mathsf{nom}(1-\prod_{j=1}^{{\ell^{\prime}}}\lambda_{d_{j}^{\prime}})<\prod_{j=1}^{\ell^{\prime}}\mathsf{denom}(\lambda_{d_{j}^{\prime}})\leq\prod\limits_{v\in V}\mathsf{denom}(\lambda_{(v,\sigma(v))})$
that
$\mathsf{common^{\prime}}\cdot\mathsf{denom}(\lambda_{e})\cdot\mathsf{denom}(w_{e})\cdot n_{2}\mbox{ is a denominator for }\mathsf{out}(\rho^{\prime})\ .$
Conequently,
$\mathsf{common^{\prime}}\cdot\mathsf{denom}(\lambda_{e})\cdot\mathsf{denom}(w_{e})\cdot n_{1}\cdot n_{2}\mbox{ is a denominator for }\mathsf{offset}(\mathsf{val}_{\sigma},e)\ .$
Putting the estimates together, we get that
$\mathsf{bound}=\mathsf{denom}(\lambda_{e})\cdot\mathsf{denom}(w_{e})\cdot\prod\limits_{v\in V}\mathsf{denom}(\lambda_{(v,\sigma(v))})^{3}\cdot\mathsf{denom}(w_{(v,\sigma(v))})$
is smaller than the smallest denominator of $\mathsf{offset}(\mathsf{val}_{\sigma},e)$. We then use this to estimate $\gamma\geq-\mathsf{offset}(\mathsf{val}_{\sigma},e)>1/\mathsf{bound}$.
@@
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be an improving DPG for a given set of positive numbers $\{\alpha_{e}>0\mid e\in E\}$, $\sigma$ a strategy for both players, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 2 for the adjusted function $f_{\sigma}^{\prime}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. Then there is a neighbouring solution $\mathsf{val}^{\prime\prime}$ to $\mathsf{val}$ such that there is a better strategy $\sigma^{\prime}$ that satisfies $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})<f_{\sigma}(\mathsf{val})$.